# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

### v1

@@ 3.1.1. Language and Toolchain Knowledge
We first evaluate whether LLMs possess the foundational knowledge required to author programs in ZK languages. To construct this evaluation, we curate a knowledge base covering the two most widely used languages, Circom and Noir. As shown in the left side of Fig. 3, sources are collected through automated web crawling of official documentation, including language specifications, tutorials, and API references (e.g., circomlib (iden3, 2025) for Circom and the Noir standard library). In addition, we manually gather publicly available examples from GitHub repositories (TEZCAN, 2025b; Sirin, [n. d.]; Contributors, [n. d.]; Thor, [n. d.]) and online tutorials (TEZCAN, 2025a) to ensure coverage of idiomatic usage.
@@
Based on this knowledge base, as shown in the middle of Fig. 3, we design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018). Inspired by earlier work on LLM-driven question generation (Liu et al., 2020; Abbasiantaeb et al., 2024; Long et al., 2024; Ji et al., 2025), our pipeline iteratively samples subsets of documents and provides them to a large model, which composes candidate questions and answers. This process is repeated until the full document collection has been exhaustively covered, producing a diverse set of drafts.
@@
All generated questions are then subjected to rigorous expert validation (right side of Fig. 3). Human reviewers examine each question for factual correctness, clarity, and alignment with the intended level of difficulty. Cross-checking is performed among the authors to ensure agreement, and questions that do not meet these standards are either corrected or discarded. Topic coverage is also verified to ensure that all major aspects of the knowledge base are represented. We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library templates and functions), and compiler principles (translation from source code to arithmetic circuits). A summary of the questions and representative examples is shown in Fig. 4. As shown, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories. The full set of questions is provided in the Artifact.
Figure 4. Illustrative MCQ on Basic Circom Syntax with a Summary of Knowledge Categories.
Figure 5. The Design of Gadget Benchmark on LLM’s Ability to Use or Implement Gadgets to Encode Constraints.
3.1.2. Using and Implementing Gadgets to Encode Constraints
The second dimension of ZK-Eval assesses the ability of LLMs to use and implement gadgets to encode constraints, which are the fundamental building blocks of ZK programs. To construct this evaluation, we curate a representative set of gadgets from three complementary sources, as shown in Fig. 5. First, circomlib (iden3, 2025), the official standard library of Circom, provides canonical implementations of common gadgets that developers routinely instantiate. Second, zk-kit (Nico Serrano, [n. d.]) extends this idea to a broader ecosystem, offering more reusable gadget implementations across multiple ZK languages. Third, as an important observation, many gadgets share similar logical and arithmetic structures with primitives in SMT solvers. To capture this perspective, we leverage the Z3Py (Research, [n. d.]) interface to the Z3 SMT solver (De Moura and Bjørner, 2008). These sources span both language-specific and language-agnostic perspectives, ensuring that our benchmark captures the core constraint patterns encountered in practice.
@@ 3.1.2. Using and Implementing Gadgets to Encode Constraints
The resulting gadget set consists of 35 gadgets covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, including addition, subtraction, multiplication, division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum, distinct, and multiplexors (see the full list in the Supplementary Material). This collection is sufficiently expressive to simulate the primitive layer of typical ZK development workflows: any non-trivial circuit can be realized through the composition of these gadgets.
For every gadget in the set, we author a concise task description. Test cases are crafted to validate both completeness: accepting input assignments and soundness: rejecting assignments. For logical gates, this corresponds to enumerating the truth table; for arithmetic and composite gadgets, it is manually constructed based on their semantics. These tests ensure that correctness is evaluated not only in terms of producing the right outputs, but also in enforcing the intended constraints.
Figure 6. The Design of End-to-end ZK Program Generation Benchmark Adapted from HumanEval.
@@ 3.1.3. End-to-End ZK Program Generation
The third dimension of ZK-Eval evaluates the capability of LLMs to perform full end-to-end ZK program generation, which requires the model to integrate language proficiency and gadget-level competence to assemble complete programs that encode non-trivial problem specifications. We build the benchmark on the top of the widely used HumanEval dataset (Chen et al., 2021) with ZKP-specific adaptions, as shown in Fig. 6. In particular, we exclude tasks involving (i) string manipulation, (ii) floating-point arithmetic, and (iii) outputs of variable length. Strings and floating-point numbers are not natively supported by Circom or Noir, while variable-length outputs are incompatible with the static nature of arithmetic circuits. Following prior work (Myrzakhan et al., 2024), each task is automatically annotated by an LLM and verified by experts to ensure accurate filtering, yielding a set of 68 standalone tasks suitable for ZK program generation.
To reflect the verification nature of ZK programs, we further reformulate the problem statements. Original HumanEval tasks are expressed as construction problems, in which a function computes an output $S=P(I)$ for input $I$. In contrast, ZK programs implement verification problems, where the goal is to check whether a candidate solution $S$ satisfies the intended relation for $I$. We therefore restate each task as a verification function $V(I,S)\in\{0,1\}$, which outputs one if and only if the candidate solution is correct.
For each task, we construct an executable Python oracle that implements the verification semantics independently of any ZK framework, serving as a ground truth reference. Following prior work (Yuan et al., 2023; Siddiq et al., 2023), we then instruct an LLM to generate two Python scripts for test case generation: one producing input-solution pairs $(I,S)$ that should be accepted, and the other producing pairs that should be rejected. This dual design is crucial, as existing code-generation benchmarks provide only accepting examples (verifying acceptance of correct input-solution pairs) but do not evaluate rejection behavior. This ensures that generated programs are tested both for soundness (rejecting invalid ones) and completeness (accepting valid ones).
To ensure the benchmark offers a rigorous and faithful assessment, all test cases are first executed against the Python oracle and then validated and cross-checked by authors for clarity, correctness, and corner-case coverage, following practices in prior benchmarks (Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022a). Whenever inconsistencies are detected at any stage, the corresponding step is re-executed or manually corrected. If corner-case is found lacking, the test case generator is revised accordingly.
@@ 3.2. Measurement Settings
  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard libraries of Circom and Noir?
  • RQ2: Gadget-Level Competence. Can LLMs reliably use or implement constraint gadgets, the building blocks of ZK programs?
  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs?
Each RQ corresponds directly to one component of ZK-Eval. For RQ1, models are tested on curated multiple-choice questions covering language knowledge. For RQ2, they are instructed to implement gadget tasks which are then automatically validated against accepting and rejecting test cases. For RQ3, they attempt end-to-end program generation from reformulated HumanEval tasks, with correctness judged by soundness and completeness against dual test suites.
Model Selection. We evaluate 4 representative LLMs: OpenAI’s reasoning models GPT-o4-mini and GPT-o3, along with DeepSeek-V3 and Qwen3. This mix of reasoning and open-source non-reasoning ones allows us to assess both family-specific and general trends. All models are run under identical prompts and budgets, with 10 samples per task to estimate pass rates and reduce variance.
@@
Figure 9. Average Error Distribution of Gadget Implementations Across Languages, Types, and Causes.
@@
Fig. 3.2 summarizes performance on the multiple-choice question set. Reasoning models perform best, with GPT-o4-mini reaching 88.1% and GPT-o3 87.2%. Open-source non-reasoning models perform noticeably worse, with DeepSeek V3 at 79.5% and Qwen3 at 78.9%. The human expert for reference scores 88.7%, similar to the reasoning models. Overall, LLMs demonstrate strong competence in ZK language knowledge, achieving accuracy within a few percentage points of the human expert.
Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-source ones and perform similar to human experts.
We further examine performance across question categories and languages. LLMs achieve their highest accuracy on questions about basic syntax and API references, reflecting familiarity with fundamental constructs and widely used library templates. In contrast, accuracy drops on advanced topics such as traits, signal taggings, and generics, which extend expressiveness but are less commonly encountered. The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits. These results indicate that while LLMs are proficient in surface-level language usage, they struggle with deeper aspects of language design and compilation. Notably, there is no significant difference between Circom and Noir, suggesting that model competence reflects general exposure to ZK languages rather than language-specific familiarity.
@@
These findings are consistent with the nature of the knowledge base: official documents, standard libraries, public tutorials, and open-source examples are all widely accessible training material to LLMs. Therefore, it is anticipated that LLMs can achieve strong performance on such content.
3.4. (RQ2) Gadget-Level Competence
@@
Building on the language knowledge observed in RQ1, we next assess whether LLMs can effectively apply this knowledge in ZK programming, specifically in using or implementing gadgets. Model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes. From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) semantic errors, where code compiled but failed functional correctness tests.
@@
Fig. 9 shows the results on the gadget benchmark. Performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir. Comparison gadgets scored lower (37% and 20%), while arithmetic and composite gadgets dropped to extremely low accuracy. This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into gadgets, LLMs largely fail to produce correct, testable implementations. This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by limited and fragmented resources (Sheybani et al., 2025).
@@
Finding ③: On Circom and Noir languages, LLMs perform consistently weak on using or implementing constraint gadgets across all categories.
@@
Typing and resolution errors are the most frequent. Noir enforces strict type matching in binary operations, disallowing implicit casts between fields and integers, while Circom distinguishes sharply between variables and signals. Models frequently mix types, reference undeclared identifiers, or redefine variables, causing compilation failures. Signal handling errors, unique to Circom, are the next most common, including declaring signals inside control-flow, reassigning them multiple times, or using them before initialization. Constraint enforcement errors also recur: instead of encoding equalities with === or assert, models generate invalid forms such as non-quadratic constraints in Circom or deprecated constructs in Noir. These issues stem from ZK-specific semantics absent in conventional languages.
@@
Performance further varies by gadget type, reflecting library support and algebraic complexity. Logical operations and comparisons, being boolean primitives or supported by libraries, are most tractable. Arithmetic and composite gadgets, such as exponentiation, modulo, floor division, or composites like sum, product, and distinct, lack native support and must be implemented from scratch. These tasks demand reasoning over field arithmetic, careful handling of inverses and negatives, and explicit soundness constraints. The very low accuracy here underscores the inability of current models to generate gadgets beyond those directly provided in libraries.
@@
Finding ④: LLMs handle boolean and comparison gadgets better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
@@
Fig. 8 summarizes the results on the end-to-end generation benchmark. Compared to language knowledge and gadget-level tasks, performance drops substantially across all models. For Circom, accuracy is very low: GPT-o3 reaches only 20% semantically correct and 22% syntactically valid programs, while Qwen3 drops to 5% and 9%. Noir results are somewhat higher but still limited, with GPT-o4-mini at 32% and 40% and Qwen3 at 10% and 19%. Across all settings, semantic correctness lags behind syntactic validity, showing that many compilable programs still fail to encode the intended verification relation.
Several trends emerge from these results. First, Noir tasks consistently yield higher success rates than Circom, likely due to Noir’s simpler type system and cleaner syntax, which reduce opportunities for errors such as signal misuse or type mismatches that are pervasive in Circom. Second, reasoning models substantially outperform open-source ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek V3 and Qwen3 across both languages, confirming the gap already observed in earlier tasks. Finally, the gap between syntactic and semantic correctness underscores that compilation alone is insufficient for assessing ZK program quality: many outputs are compilable but wrong.
Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-source ones; Noir is easier than Circom; many programs compile but remain incorrect.

### v2

@@ 3.1.1. Language and Toolchain Knowledge
We first evaluate whether LLMs possess the foundational knowledge required to author ZK programs. As shown in the left side of Fig. 3, sources are collected through web crawling of official documentation, including language specifications, tutorials, and API references (e.g., Circom’s circomlib (iden3, 2025) and the Noir’s stdlib). Additionally, we gather publicly available pedagogical examples from GitHub repositories (TEZCAN, 2025b; Sirin, ; Contributors, ; Thor, ) and online tutorials (TEZCAN, 2025a) to increase coverage of idiomatic usage. This yields 51 documents for Circom and 40 for Noir.
@@
Based on this knowledge base, we then design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018). Inspired by earlier work on LLM-driven question generation (Liu et al., 2020; Abbasiantaeb et al., 2024; Long et al., 2024; Ji et al., 2025), our pipeline iteratively samples subsets of documents and provides them to a large model, which composes candidate questions and answers. This process is repeated until the full document collection has been exhaustively covered, producing a diverse set of drafts.
@@
All generated questions are then subjected to rigorous expert validation. Three human reviewers independently assess each item for factual correctness, clarity, and alignment with the intended difficulty, with disagreements resolved through discussion to consensus. Cross-checking by the authors ensures consistency across the set. Topic coverage is also verified to ensure that all major aspects (as listed in the official documents) of the studied language are represented. We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library APIs), and compiler principles (translation from source code to arithmetic circuits). As shown in Fig. 3, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories. Representative examples are provided in the Supplementary Material and the full set of questions is provided in the Artifact.
3.1.2. Implementing and Using Algebraic Primitives to Specify Constraints
We assess the capability of LLMs to implement and utilize algebraic primitives as the building blocks of ZK constraint systems. Different from higher-level cryptographic (e.g., hash functions, signatures) and application-level (e.g., account-balance checks, order-matching validity) APIs available in external libraries, these encapsulate the core relational logic that developers must wire to define custom circuit behavior.
Technical Difficulty. We clarify that, constructing a benchmark that reflects the fundamental algebraic patterns essential to real-world ZK programs is challenging. In existing codebases, core relational logic is often entangled with application-specific boilerplate, making it difficult to isolate and evaluate the underlying algebraic reasoning. To decouple and isolate these foundational patterns, we systematically curate a set of primitives from widely adopted libraries. This process, illustrated in Fig. 3, draws from three complementary sources: circomlib (iden3, 2025), zk-kit (Nico Serrano, ), and Z3Py (Research, ).
  1. (1)
    Standardized Libraries. We first include canonical primitives from circomlib and zk-kit, which represent the production standard for routine circuit instantiation.
  2. (2)
    Agnostic Extraction. From these repositories, we manually extract and deduplicate algebraic primitives while filtering out application-specific composites. This ensures the benchmark captures the universal building blocks of ZK logic rather than application-specific patterns.
  3. (3)
    Aligned Completeness. To ensure comprehensiveness, we cross-reference our set with the Z3Py SMT-solver interface and incorporate primitives with similar structures, guaranteeing completeness for algebraic constraint specification.
@@ 3.1.2. Implementing and Using Algebraic Primitives to Specify Constraints
The resulting benchmark consists of 35 algebraic primitives covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, such as addition, multiplication, integer division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum and multiplexers (see the full list in the Supplementary Material). This collection serves as a complete basis for constraint specification; as these algebraic primitives represent the atomic operations of arithmetic circuits, any non-trivial verification relation can be realized through their rigorous composition.
Practical Relevance. To check the practical relevance of our benchmark, we surveyed 37 production-grade ZK repositories comprising 408 source files. Our analysis confirms that the selected algebraic primitives are highly representative of real-world circuit logic: we achieve 97.30% and 94.59% coverage for core arithmetic and relational primitives, respectively, which are universal to production ZK systems. While logical (70.27%) and specialized composites (16.22%) appear less frequently, they remain mathematically essential for logical completeness and the construction of complex verification relations. Notably, it includes every one of the top 10 most frequently used primitives identified in the survey, establishing a representative foundation.
Task Setup. For each primitive, we author a concise task description that asks the LLM to implement the primitive’s functional logic and use it to specify the intended algebraic relation. Following established practice on test-case curation (Chen et al., 2021; Jain et al., 2024), we craft test-cases that validate both completeness (accepting valid witnesses) and soundness (rejecting invalid assignments). For finite primitives (e.g., logical xor), test-cases exhaustively enumerate the entire truth table. For primitives with larger domains, we combine automated random input generation with edge-cases (e.g., field overflows, zero-value or out-of-domain inputs), yielding at least 100 cases per primitive. To validate test rigor, we injected artificial errors into ground-truth implementations (like a mutation testing process); our tests detected 87.04% of these faults, confirming their effectiveness at catching subtle mistakes.
@@ 3.1.3. End-to-End ZK Program Generation
The final stage assesses LLMs to generate complete and correct ZK programs to verify computational statements. Building upon the first two stages, it evaluates the systemic integration of these skills to construct sound, multi-primitive constraint systems that faithfully encode high-level verification relations.
Benchmark Design Rationale. To ensure a rigorous and trustworthy assessment in the absence of a standardized ZK-specific generation dataset, we leverage canonical algorithmic baselines from the well-established HumanEval dataset. We posit that ZK programming represents a fundamental paradigm shift from imperative computation to algebraic verification rather than a shift in the application domain itself. Since the core logic of production-grade ZK applications (e.g., private finance, decentralized identity, and state-transition proofs) relies on universal algorithmic patterns (e.g., sorting, membership checks, and conditional branching), adapting these general algorithmic tasks allows us to isolate the gap on computation-to-verification translation without the confounding variables of application-specific requirements (e.g., application-specific API calling).
To reflect the unique semantics of ZK proofs, we systematically transform the original benchmark through three distinct adaptations, as illustrated in the right column of Fig. 3:
  1. (1)
    Compatibility Filtering. We exclude tasks involving string manipulation and floating-point arithmetic which are not supported by DSLs. Variable-length output tasks are also excluded since they’re incompatible with the static nature of arithmetic circuits. Their removal ensures that evaluation focuses on the model’s constraint-reasoning ability rather than DSL limitations. This yields a curated set of 68 standalone tasks suitable for ZK synthesis.
  2. (2)
    Verification Reformulation. We reformulate the natural language description of each computational task (defined as finding $S=P(I)$ for input $I$) into a verification problem. The LLM must generate a verification function $V(I,S)\in\{0,1\}$ that returns $1$ if and only if the candidate solution $S$ satisfies the intended relation for input $I$.
  3. (3)
    Oracle Adaptation. Correspondingly, the ground-truth oracle solutions are rewritten from computational functions ($S=P(I)$) to verification forms ($V(I,S)$). To ensure these serve as a reliable ground truth, we adopt a declarative implementation style; for instance, a sorting task is verified by checking permutation and monotonicity predicates rather than re-executing a sorting algorithm. To align with the semantics of ZKPs, these oracles are restricted to finite field arithmetic (mod $p$) and avoid non-deterministic execution paths. Each oracle was manually authored and cross-validated against the original computational benchmarks to ensure its reliability as a reference for test-case preparation.
Quality Assurance. Following established practices in benchmark construction (Yuan et al., 2023; Siddiq et al., 2023; Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022), we implement a rigorous validation process to ensure the Soundness and Completeness of generated programs. While traditional datasets provide only “Acceptance Cases” (Completeness), ZK programs require “Rejection Cases” to ensure the circuit is properly constrained (Soundness). Thus we implement test-case generators to produce input-solution pairs that a correct verifier must reject, ensuring soundness validation. We cross-validate task descriptions, adapted oracles, and test-case generators through manual review and injection of edge cases (e.g., all-zeros, field-overflows, out-of-domain inputs, etc.). Collectively, this process yields 100 test cases per task with an average distribution of 54.41% acceptance and 45.59% rejection cases. To quantify benchmark quality, we injected syntactic mutations into the ground-truth oracles (mutation-testing); our test suites detected 86.42% mutations, demonstrating high sensitivity to subtle errors.
@@ 3.2. Measurement Settings
  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard library APIs of Circom and Noir?
  • RQ2: Algebraic Primitive Competence. Can LLMs reliably implement and use algebraic primitives, the building blocks of ZK programs, to specify correct constraints?
  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs on the intended computational relation?
Each RQ corresponds directly to one component of ZK-Eval. For RQ1, models are tested on curated multiple-choice questions covering language knowledge. For RQ2, they are instructed to implement and use algebraic primitives which are then automatically validated against acceptance and rejection test cases. For RQ3, they attempt end-to-end program generation to realize verification constraints for general computational problems, whose correctness is judged by soundness and completeness against dual test suites.
Model Selection. We evaluate 4 representative LLMs: OpenAI’s reasoning models GPT-o4-mini and GPT-o3, along with DeepSeek-V3 and Qwen3. This mix of reasoning and open-weight non-reasoning ones allows us to assess both family-specific and general trends. All models are run under identical prompts and budgets, with 10 samples per task to estimate Pass@1 rates and reduce variance.
@@
Figure 6. Average Error Distribution of Primitive Implementations Across Languages, Types, and Causes.
@@
Fig. 3.2 summarizes performance. Reasoning models perform best, with GPT-o4-mini reaching 88.1% and GPT-o3 87.2%. Open-weight non-reasoning models perform noticeably worse, with DeepSeek-V3 at 79.5% and Qwen3 at 78.9%. The human expert for reference scores 88.7%, similar to the reasoning models. Overall, LLMs demonstrate strong competence in ZK language knowledge, achieving accuracy within a few percentage points of the human expert.
Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-weight ones and perform similar to human experts.
We further examine performance across question categories and languages. LLMs achieve their highest accuracy on questions about basic syntax and API references, reflecting familiarity with fundamental constructs and widely used library templates. In contrast, accuracy drops on advanced topics such as traits, signal taggings, and generics, which extend expressiveness but are less commonly encountered. The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits, indicating that they struggle with deeper aspects of arithmetic circuits and constraint semantics on finite fields. Notably, there is no significant difference between Circom and Noir, suggesting that model competence reflects general exposure to ZK languages rather than language-specific familiarity.
@@
These findings are consistent with the nature of the knowledge base: official documents, public tutorials, and examples are all widely accessible training material to LLMs. Thus, it is anticipated that LLMs can achieve strong performance on such content. This suggests that (i) explicit augmentations on language grammars or official documentations may be unnecessary (ii) LLMs exhibit strong potential for ZK programming tasks but require further evaluations to reveal the bottlenecks.
3.4. (RQ2) Algebraic Primitive Competence
@@
Language knowledge enables syntactic fluency, but ZK programming fundamentally requires the ability to specify algebraic constraints. We next assess whether LLMs can use or implement algebraic primitives, the atomic elements of constraint systems. Specifically, model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes. From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) functional errors, where code compiled but failed functional correctness tests.
@@
Fig. 6 shows the results and performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir. Comparison scored lower (37% and 20%), while Arithmetic and Composites dropped to extremely low accuracy. This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into specifying constraints with algebraic primitives, LLMs largely fail to produce testable, correct implementations. This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by the inherent programming paradigm shift (Sheybani et al., 2025).
@@
Finding ③: On Circom and Noir languages, LLMs perform consistently weak on implementing and using algebraic primitives to specify constraints across all categories.
@@
Typing & resolution errors are the most frequent. Noir enforces strict type matching, disallowing implicit casts between fields and integers, while Circom distinguishes between variables and signals. Models frequently mix types, reference undeclared identifiers, or redefine variables, causing compilation failures. Signal handling errors, unique to Circom, are the next most common, including declaring signals inside control-flow, reassigning them multiple times, or using them before initialization. Constraint enforcement errors also recur: models generate invalid constraint forms such as non-quadratic constraints (e.g., x * y * z === 1) or use off-circuit assertions. These issues stem from ZK-specific semantics absent in conventional languages.
@@
Performance further varies by primitive type, reflecting library support and algebraic complexity. Logical operations and comparisons, being boolean primitives or supported by libraries, are most tractable. Arithmetic and Composites, such as exponentiation, modulo, floor division, or composites like sum and product, lack native support and must be implemented from scratch. These tasks demand reasoning over field arithmetic, careful handling of inverses and negatives, and explicit soundness constraints. The very low accuracy here underscores the inability of current models to generate implementations beyond those directly provided in libraries.
@@
Finding ④: LLMs handle boolean and comparison primitives better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
@@
Fig. 5 summarizes the results on the end-to-end generation benchmark. Compared to language knowledge and primitive-level tasks, performance drops substantially across all models. For Circom, accuracy is very low: GPT-o3 reaches only 20% semantically correct and 22% syntactically valid programs, while Qwen3 drops to 5% and 9%. Noir results are somewhat higher but still limited, with GPT-o4-mini at 32% and 40% and Qwen3 at 10% and 19%.
Several trends emerge from these results. First, Noir tasks consistently yield higher success rates than Circom, likely due to Noir’s simpler type system and cleaner syntax, which reduce opportunities for errors such as signal misuse or type mismatches that are pervasive in Circom. Second, reasoning models substantially outperform open-weight ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek-V3 and Qwen3 across both languages, confirming the model capability difference already observed in earlier tasks. Finally, the gap between syntactic and semantic correctness underscores that being compilable alone is insufficient for assessing ZK program quality: many outputs are compilable but suffer from under- or over-constrained relations.
Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-weight ones; Noir is easier than Circom; many programs compile but remain incorrect.
This reveals the fundamental challenge of ZK programming: it requires not only language fluency and primitive competence, but also the ability to architect multi-primitive constraint systems that faithfully encode complex verification relations — a skill that current LLMs have not yet mastered.